# BillBoard Splatting (BBSplat): Learnable Textured Primitives for Novel View Synthesis

### v3

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. The proposed primitives close the rendering quality gap between 2D and 3D Gaussian Splatting (GS), preserving the accurate mesh extraction ability of 2D primitives. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in the storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. Project page: david-svitov.github.io/BBSplat_project_page.
@@ 1 Introduction
Recently, 2D Gaussian Splatting (2DGS) [20] proposed the use of flat Gaussians, oriented in 3D as flat primitives align better with the object surface. Since 2D Gaussians are effectively tangent to object surfaces, they allow more accurate surface extraction. Despite proving their efficiency in mesh extraction task, 2D primitives result in a downgrade of rendering metrics compared to 3D primitives such as 3D Gaussians. In this work, we aim to make 2D primitives suitable for high-quality NVS by intoducing a new primitive representation.
@@
  • The BBSplat representation allows extracting accurate meshes from a scene with accuracy superior to 3D Gaussian-based methods.
  • We developed an algorithm to efficiently represent and store textures for billboards. BBSplat demonstrates significant storage cost reduction. As a result, for some scenarios, storing the scene represented with billboards is more efficient than 3DGS or 2DGS.
@@ 2 Related work
Recent works further improve 3DGS in terms of rendering quality and speed. Thus, in [6] authors revise the densification algorithm to reduce the number of artifacts by using loss-based splitting of Gaussians. In 3DGS-MCMC [25], authors use Markov Chain Monte Carlo interpretation of sampling algorithm to more efficiently distribute Gaussians over the scene, even with random initialization. RadSplat [37] uses pre-trained NeRF representation as a prior during 3DGS training to improve rendering quality in challenging scenarios. Mip-Splatting [53] solves the issue of gaps between Gaussian produced due to changes in camera distance by introducing a 3D smoothing filter that constrains Gaussian size.
@@ 3.1 Preliminaries: 2D Gaussian Splatting
where matrices $R_{i}$ and $S_{i}$ produced from quaternion rotation $r_{i}$ and scale $s_{i}$, and $(u,v)$ is the point coordinate on the plane $\mathcal{P}_{i}$. Then, 2DGS utilizes $(u,v)$ coordinates of points within a plane to calculate 2D Gaussian intensity as follows:
@@ 3.2 Billboard Splatting
In this work, we propose to further improve NVS efficiency and quality by introducing billboard splatting. These billboard primitives are based on planes, but leverage learnable textures instead of utilizing Gaussian distribution (Eq. 2) to calculate their color and transparency. We inherit 2DGS parametrization for these planes: $\{\mu_{i},s_{i},r_{i},\textrm{SH}_{i}\}$, where $\mu_{i}$ is the center of $i$-th plane, $s_{i}$ is scales along two axis, $r_{i}$ is rotation quaternion, $\textrm{SH}_{i}$ is spherical harmonics coefficients. Instead of using global Gaussian opacity $o_{i}$ we set transparency at each point of a plane as $T_{i}^{\alpha}$ texture (Fig. 1 (b2)). This way, each billboard can have an arbitrary shape. Additionally, we parametrize billboards with $T_{i}^{\textrm{RGB}}$ RGB texture to control the color of all points of the billboard.
We use explicit ray-splat intersection algorithm [43] to find the corresponding plane coordinate $\textbf{u}=(u,v)$ for $\textbf{x}=(x,y)$ screen coordinate. Specifically, we parametrize the camera ray as an intersection of two 4D homogeneous planes: $h_{x}=(-1,0,0,x)^{T}$ and $h_{y}=(0,-1,0,y)^{T}$. We then transform these planes to the $uv$-coordinate system of plane $H$ to find intersection point in $(u,v)$ coordinates:
$h_{u}=(WH)^{T}h_{x}\hskip 28.45274pth_{v}=(WH)^{T}h_{y},$ (3)
@@
$u(x)=\dfrac{h_{u}^{2}h_{v}^{4}-h_{u}^{4}h_{v}^{2}}{h_{u}^{1}h_{v}^{2}-h_{u}^{2}h_{v}^{1}}\hskip 14.22636ptv(x)=\dfrac{h_{u}^{4}h_{v}^{1}-h_{u}^{1}h_{v}^{4}}{h_{u}^{1}h_{v}^{2}-h_{u}^{2}h_{v}^{1}},$ (4)
where $h_{u}^{i},h_{v}^{i}$ are the $i$-th value of the 4D homogeneous plane parameters.
@@
To calculate value of view-dependent colors $c_{i}$ at each uv-point u of $i$-th billboard, we integrate sampled texture colors $T_{i}^{\textrm{RGB}}[\textbf{u}]$ as offsets to the colors predicted with $\textrm{SH}_{i}$ and view-direction vector $\vec{d_{i}}$:
$c_{i}[\textbf{u}]=T_{i}^{\textrm{RGB}}[\textbf{u}]+RGB(\textrm{SH}_{i},\vec{d_{i}}).$ (6)
Here $\textrm{SH}_{i}$ is a spherical harmonic features [2] for each primitive, and $\vec{d_{i}}$ is the view-direction vector from the camera position to billboard center position $\mu_{i}$. We transform $\textrm{SH}_{i}$ to RGB color space and define this operation as $RGB(\cdot,\cdot)$. This way, we can handle view-dependent light effects for textured planes (Fig. 2). Representing the final color as a combination of spherical harmonics and sampled color was initially proposed in [50] and helps to disentangle texture and light effects. In our case it also allows $T_{i}^{\textrm{RGB}}$ to have a sparse structure, which is more efficient to store as described in Section 3.5.
@@ 3.3 Training
Figure 2: Texture rasterization. During rendering to get final billboard colors we add texture $T^{\textrm{RGB}}_{i}$ colors to base color calculated with $\textrm{SH}_{i}$ and view-direction vector $\vec{d_{i}}$. It results in textured billboards handling light effects. Here we showcase four possible renderings for different camera directions.
@@ 3.4 Adaptive density control
This renders parameters adjustment more straightforward compared to Eq. 12. Finally, where MCMC compares opacity $o_{i}$ with a predefined threshold $\gamma=5\mathrm{e}{-3}$ to determine “dead” Gaussians, we compare the average value of the alpha-map: $\overline{T_{i}^{\alpha}}<\gamma$.
@@ 3.5 Texture compression
Figure 3: Tanks&Temples metrics. We report quantitative metrics values depending on the number of primitives for each method. BBSplat outperforms state-of-the-art for the lower number of primitives.
Our final color representation described in Eq. 6 as a sum of color calculated with SH and texture sampled offsets allows us to additionally reduce storage costs. Our regularization term $\mathcal{L_{\textrm{texture}}}$ proposed in Eq. 9 allows us to get sparser $T^{\textrm{RGB}}$ values and $T^{\alpha}$ closer to a Gaussian distribution $\mathcal{G}$. We increase the sparsity of $T^{\alpha}$ by subtracting the Gaussian pattern $T_{\alpha}-\mathcal{G}$ and ensure more zero values in the texture. The sparse nature of textures allows us to use efficient dictionary-based compression methods [42, 11] (e.g. ZIP [1]) to further reduce storage costs. Our texture compression pipeline allows us to reduce memory consumption $\sim 7$ times on average.
@@ 4 Experiments
Table 1: Quantitative results with the same number of primitives. We report metrics for the following datasets: Tanks&Temples [27], Mip-NeRF-360 [5], and DTU [23]. We fixed a number of primitives for all methods to provide metrics for the same training scenario. †- denotes the method’s version without using depth-normal regularization.
@@
Table 2: Quantitative results with a large number of primitives. We report metrics for the following datasets: Tanks&Temples [27], Mip-NeRF-360 [5], and DTU [23]. We provide metrics for the recommended number of Gaussians for each method, which results in increased storage space.†- indicates method version without using depth-normal regularizations.
Figure 4: Qualitative results. We provide rendering results of scenes from each dataset: Tanks&Temples [Train], Mip-NeRF-360 [Room], and DTU [Scan97]. For competitors, we use the maximum recommended by the method number of Gaussians. Our results are reported for 300K billboards. More renderings can be found in the Sec. 7.1 of supp. mat..
@@
Texture-GS 30.03 0.0135 0.1440
Ours-small 30.81 0.0123 0.1055
@@ 4.1 Implementation
For training, we use the Adam [26] optimizer with hyper-parameters recommended for the Gaussian splatting pipelines [24]. We set RGB and $\alpha$ textures size $S_{T}=16$. resulting in $16\times 16$ textures to get the best trade-off between quality and memory requirements. To optimize RGB-texture we set learning rate $lr_{\textrm{RGB}}=2.5\mathrm{e}{-3}$ and for $\alpha$-texture we set $lr_{\alpha}=1\mathrm{e}{-3}$. Additionally, we found it beneficial for convergence to increase spherical harmonics learning rate to $lr_{\textrm{SH}}=5\mathrm{e}{-3}$.
We implemented texture sampling with back-propagation in the CUDA code to achieve fast inference.
@@ 4.2 Results and Evaluation
We compared with previous state-of-the-art methods on real-world scenes using standard datasets. We evaluated our method on 4 indoor scenes from the Mip-NeRF-360 dataset [5] and 5 outdoor scenes from the Tanks&Temples dataset [27]. Additionally, we evaluated our method on 15 scenes from DTU dataset [23]. Overall, we provide experiments for 24 real scenes from three datasets.
@@
We used train/test split for all datasets following 3DGS methodology where each 8th image was taken for the test. For all three datasets, we used images downsampled to 800 pixels on the larger side. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
@@
Quantitative comparison. We report PSNR, SSIM, and LPIPS [56] metrics (Tabs. 1 and 2) widely used to evaluate quality in the NVS task. PSNR and SSIM are traditionally used to demonstrate similarity with ground truth images; LPIPS utilize neural networks and is known for the best correlation with human perception. In the provided experiments, we evaluated metrics for different numbers of primitives by limiting the growth of their number during training as proposed in [6].
@@
In an experiment in Table 1, we showcase metrics for the fixed number of primitives depending only on their type. For the same amount of primitives, our method reaches better PSNR, SSIM and LPIPS than state-of-the-art Gaussian-based representations both for indoor and outdoor scenes. While improving state-of-the-art metrics, for the same number of primitives BBSplat requires more storage space.
In Table 2, we compare against other methods for the maximum recommended amount of Gaussians to achieve the best possible quality for each method. For our method, we provide two scenarios with fewer (small) and more (big) billboards.
In Figure 3, we show the metrics as a function of the number of primitives (Gaussians or BBSplats). To plot the leftmost point we used the maximum number of Gaussians requested by each method, then for subsequent points we fixed the number of primitives for all scenes at the chosen value. It can be seen that our method gives a flatter curve of metrics change, which means a minimal decrease in metrics when the number of primitives decreases to 80K.
@@
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to Sec. 7.1 of supp. mat..
@@
Mesh Extraction. In Table 4, we compare the 3D mesh extraction performance of BBSPlat with state-of-the-art methods on the DTU dataset [23]. For the experiment, we use 15 scenes with ground truth obtained by 3D scanning objects in a laboratory setting. As a metric, we present the Chamfer distance indicating the similarity between two meshes. Both experiments for 2DGS and for BBSplat were conducted using the depth-normal regularization proposed in [20]. Rendering quality metrics for BBSplat in Tables 1 and 2 were also obtained using such regularizations. BBSplat reaches 0.91 Chamfer distance on DTU dataset, which is better than 1.96 for 3DGS and 1.33 for SuGaR, and close to 0.80 for 2DGS.
Method Chamfer distance
3DGS [24] 1.96
SuGaR [17] 1.33
2DGS [20] 0.80
BBSplat(Ours) 0.91

### v4

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for novel view synthesis based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. The proposed primitives close the rendering quality gap between 2D and 3D Gaussian Splatting (GS), enabling the accurate extraction of 3D mesh as in the 2DGS framework. Additionally, the explicit nature of planar primitives enables the use of the ray-tracing effects in rasterization. Our novel regularization term encourages textures to have a sparser structure, enabling an efficient compression that leads to a reduction in the storage space of the model up to ×$\times 17$ times compared to 3DGS. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. Namely, we achieve a state-of-the-art PSNR of 29.72 for DTU at Full HD resolution. Project page: david-svitov.github.io/BBSplat_project_page.
@@ 1 Introduction
Recently, 2D Gaussian Splatting (2DGS) [20] proposed the use of flat Gaussians, oriented in 3D as flat primitives align better with the object surface. Since 2D Gaussians are effectively tangent to object surfaces, they allow more accurate surface extraction. Despite proving their efficiency in mesh extraction tasks, 2D primitives result in a downgrade of rendering metrics compared to 3D primitives such as 3D Gaussians. In this work, we aim to make 2D primitives suitable for high-quality NVS by introducing a new primitive representation, but still allowing reliable mesh extractions.
@@
  • The BBSplat representation allows extracting reliable meshes from a scene with accuracy superior to 3D Gaussian-based methods.
  • We developed an algorithm to efficiently represent and store textures for billboards. BBSplat demonstrates significant storage cost reduction.
@@ 2 Related work
Recent works further improve 3DGS in terms of rendering quality. Thus, in [6] authors revise the densification algorithm to reduce the number of artifacts by using loss-based splitting of Gaussians. In 3DGS-MCMC [25], authors use Markov Chain Monte Carlo interpretation of sampling algorithm to more efficiently distribute Gaussians over the scene, even with random initialization. RadSplat [38] uses pre-trained NeRF representation as a prior during 3DGS training to improve rendering quality in challenging scenarios. Mip-Splatting [56] solves the issue of gaps between Gaussian produced due to changes in camera distance by introducing a 3D smoothing filter that constrains Gaussian size.
@@ 3.1 Preliminaries: 2D Gaussian Splatting
where matrices $R_{i}$ and $S_{i}$ are given by quaternion rotation $r_{i}$ and scale $s_{i}$, and $(u,v)$ is the point coordinate on the plane $\mathcal{P}_{i}$. Then, 2DGS utilizes $(u,v)$ coordinates of points within a plane to calculate 2D Gaussian intensity as follows:
@@ 3.2 Billboard Splatting
In this work, we propose to further improve NVS efficiency and quality by introducing billboard splatting. These billboard primitives are based on planes, but leverage learnable textures instead of utilizing Gaussian distribution (Eq. 2) to calculate their color and transparency. We inherit 2DGS parametrization for these planes: $\{\mu_{i},s_{i},r_{i},\textrm{SH}_{i}\}$, where $\mu_{i}$ is the center of $i$-th plane, $s_{i}$ is scales along two axis, $r_{i}$ is rotation quaternion, $\textrm{SH}_{i}$ is spherical harmonics coefficients. Instead of using global Gaussian opacity $o_{i}$ we set transparency at each point of a plane as $T_{i}^{\alpha}$ texture (Fig. 1 (b2)). This way, each billboard can have an arbitrary shape. Additionally, we parametrize billboards with $T_{i}^{\textrm{RGB}}$ a RGB texture to control the color of all points of the billboard.
We use explicit ray-splat intersection algorithm [46] to find the corresponding plane coordinate $\textbf{u}=(u,v)^{T}$ for $\textbf{x}=(x,y)^{T}$ screen coordinate. Specifically, we parametrize the camera ray as an intersection of two 4D homogeneous planes: $\textbf{h}_{x}=(-1,0,0,x)^{T}$ and $\textbf{h}_{y}=(0,-1,0,y)^{T}$. We then transform these planes to the $uv$-coordinate system of plane $H$ to find intersection point in $(u,v)$ coordinates:
$\textbf{h}_{u}=(WH)^{T}\textbf{h}_{x}\hskip 28.45274pt\textbf{h}_{v}=(WH)^{T}\textbf{h}_{y},$ (3)
@@
$\textbf{u}(x)=\dfrac{\textbf{h}_{u}^{2}\textbf{h}_{v}^{4}-\textbf{h}_{u}^{4}\textbf{h}_{v}^{2}}{\textbf{h}_{u}^{1}\textbf{h}_{v}^{2}-\textbf{h}_{u}^{2}\textbf{h}_{v}^{1}}\hskip 14.22636pt\textbf{v}(x)=\dfrac{\textbf{h}_{u}^{4}\textbf{h}_{v}^{1}-\textbf{h}_{u}^{1}\textbf{h}_{v}^{4}}{\textbf{h}_{u}^{1}\textbf{h}_{v}^{2}-\textbf{h}_{u}^{2}\textbf{h}_{v}^{1}},$ (4)
where $\textbf{h}_{u}^{i},\textbf{h}_{v}^{i}$ are the $i$-th value of the 4D homogeneous plane parameters.
@@
To calculate value of view-dependent colors $c_{i}$ at each uv-point u of $i$-th billboard, we integrate sampled texture colors $T_{i}^{\textrm{RGB}}[\textbf{u}]$ as offsets to the colors predicted with $\textrm{SH}_{i}$ and view-direction vector $\textbf{d}_{i}$:
$c_{i}[\textbf{u}]=T_{i}^{\textrm{RGB}}[\textbf{u}]+RGB(\textrm{SH}_{i},\textbf{d}_{i}).$ (6)
Here $\textrm{SH}_{i}$ is a spherical harmonic features [2] for each primitive, and $\textbf{d}_{i}$ is the view-direction vector from the camera position to billboard center position $\mu_{i}$. We transform $\textrm{SH}_{i}$ to RGB color space and define this operation as $RGB(\cdot,\cdot)$. This way, we can handle view-dependent light effects for textured planes (Fig. 2). Representing the final color as a combination of spherical harmonics and sampled color was initially proposed in [53] and helps to disentangle texture and light effects. In our case it also allows $T_{i}^{\textrm{RGB}}$ to have a sparse structure, which is more efficient to store as described in Section 3.5.
@@
The use of alpha maps in billboard splitting is crucial for accurate mesh extraction as it allows for detailed scene representation (Fig. 4 (b)) even for a limited number of billboards. We then render depth maps corresponding to training views and fuse them with truncated signed distance fusion (TSDF) for mesh extraction, using Open3D [61].
@@ 3.3 Training
Figure 2: Texture rasterization. During rendering to get final billboard colors we add texture $T^{\textrm{RGB}}_{i}$ colors to base color calculated with $\textrm{SH}_{i}$ and view-direction vector $\textbf{d}_{i}$. It results in textured billboards handling light effects. Here we showcase four possible renderings for different camera directions.
@@ 3.4 Adaptive density control
This makes the adjustment on parameters more straightforward compared to Eq. 12. Finally, where MCMC compares opacity $o_{i}$ with a predefined threshold $\gamma=5\mathrm{e}{-3}$ to determine “dead” Gaussians, we compare the average value of the alpha-map: $\overline{T_{i}^{\alpha}}<\gamma$.
@@ 3.5 Texture compression
Our final color representation described in Eq. 6 as a sum of color calculated with SH and texture sampled offsets allows us to additionally reduce storage costs. Our regularization term $\mathcal{L_{\textrm{texture}}}$ proposed in Eq. 9 allows us to get sparser $T^{\textrm{RGB}}$ values and $T^{\alpha}$ closer to a Gaussian distribution $\mathcal{G}$. We increase the sparsity of $T^{\alpha}$ by subtracting the Gaussian pattern $T^{\alpha}-\mathcal{G}$ and ensure more zero values in the texture. The sparse nature of textures allows us to use efficient dictionary-based compression methods [45, 11] (e.g. ZIP [1]) to further reduce storage costs. Our texture compression pipeline allows us to reduce memory consumption $\sim 7$ times on average.
Figure 3: Qualitative results. We provide rendering results of scenes from each dataset: Tanks&Temples [Train], Mip-NeRF-360 [Room], and DTU [Scan97]. For competitors methods, we use the maximum number of Gaussians recommended by the method. More renderings can be found in the Sec. 7.1 of Supp. Mat..
@@ 4 Experiments
Table 1: Quantitative results with a fixed storage space. We report metrics for the following datasets: Tanks&Temples [27], Mip-NeRF-360 [5], and DTU [23]. We demonstrate metrics for a scenario where we limit the number of primitives to achieve a small storage space. †- denotes the method’s version without using depth-normal regularization.
@@
Table 2: Quantitative results with a large number of primitives. We report metrics for the following datasets: Tanks&Temples [27], Mip-NeRF-360 [5], and DTU [23]. We provide metrics for the default number of Gaussians for each method, which results in increased storage space.†- indicates method version without using depth-normal regularizations.
Figure 4: Ray-tracing and mesh extraction. a) BBSplat representation of the scene as explicit textured planes allows the implementation of ray-tracing effects during rasterization. b) Rasterization of planes with alpha texture provide accurate objects suitable for mesh extraction. c) Alternatively, our representation can be used to extract accurate 3D meshes, surpassing the accuracy of 3D Gaussians-based methods (e.g. SuGaR). More renderings and extracted meshes can be found in the Sec. 7.1 of Supp. Mat..
@@
3DGS [24] 28.39 0.8775 0.2850 121MB
Reduced-3DGS [39] 28.31 0.8769 0.2859 30 MB
Compact-3DGS [30] 28.39 0.8752 0.3000 10MB
BBSplat(Ours) 29.39 0.8830 0.2770 7 MB
Table 3: Comparison with compression methods on DTU. We provide better NVS than state-of-the-art compression methods while requiring less storage space.
@@ 4.1 Implementation
For training, we use the Adam [26] optimizer with hyper-parameters recommended for the Gaussian splatting pipelines [24]. We set RGB and $\alpha$ textures size $S_{T}=16$. resulting in $16\times 16$ textures to get the best trade-off between quality and memory requirements. To optimize RGB-texture we set learning rate $lr_{\textrm{RGB}}=2.5\mathrm{e}{-3}$ and for $\alpha$-texture we set $lr_{\alpha}=1\mathrm{e}{-3}$. Additionally, we found it beneficial for convergence to increase spherical harmonics learning rate to $lr_{\textrm{SH}}=5\mathrm{e}{-3}$. For more details, please refer to the supplementary materials.
@@ 4.2 Results and Evaluation
We compared with previous state-of-the-art methods on real-world scenes using standard datasets. We evaluated our method on 7 publicly available scenes from the Mip-NeRF-360 dataset with indoor and outdoor real scenes [5] and also on 5 outdoor scenes from the Tanks&Temples dataset [27]. Furthermore, we evaluated our method on 15 scenes from DTU dataset [23]. Overall, we provide experiments for 27 real scenes from three datasets.
@@
We used train/test split for all datasets following 3DGS methodology where each 8th image was taken for the test. For all three datasets, we use the same image resolution as previous methods for fair comparison (e.g. 3DGS). Namely, we reduced the original images from Tanks&Temples ×$\times 2$ times (to $\sim$900px width), Mip-NeRF-360 indoor scenes ×$\times 2$ times (to $\sim$1600px width), outdoor scenes ×$\times 4$ times (to $\sim$1200px width). For DTU we used original image size ($1600\times 1200$px). All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
@@
Quantitative comparison. We report PSNR, SSIM, and LPIPS [59] metrics (Tabs. 1 and 2) widely used to evaluate quality in the NVS task. PSNR and SSIM are traditionally used to demonstrate similarity with ground truth images; LPIPS utilizes neural networks and it is known for the best correlation with human perception. In the provided experiments, we evaluated metrics for limited storage space constraining the amount of primitives by limiting the growth of their number during training as proposed in [6].
@@
In an experiment in Table 1, we showcase metrics for the fixed storage space for different NVS methods. For the same storage space, our method reaches better PSNR, SSIM and LPIPS than state-of-the-art Gaussian-based representations both for indoor DTU and outdoor Tanks&Temples scenes. For the Mip-NeRF-360 dataset, BBSplat reaches the best LPIPS even for highly detailed outdoor scenes challenging for textured representation. Notice also that BBSplat at fixed memory uses less primitives than competing methods, showing the representation power of textured Gaussians.
In Table 2, we compare against other methods for the maximum recommended amount of Gaussians to achieve the best possible quality for each method. BBSplat achieves state-of-the-art PSNR value for indoor DTU dataset while reaching up to ×$\times 5$ storage space reduction. For Tanks&Temples outdoor scenes, we provide second best NVS quality reducing storage almost ×$\times 2$ times. The Mip-NeRF-360 dataset is particularly challenging for texture-based representation as it contains many highly detailed outdoor scenes. Even in such a disadvantageous scenario, BBSplat reaches a high NVS quality. In the Supp. Mat. we provide per-scene metrics demonstrating that BBSplat provides high metrics values for indoor Mip-NeRF-360 scenes and lower metrics for outdoor scenes, which is one of the method’s limitations.
In Table 3, we provide a comparison with state-of-the-art compression methods for 3DGS. All provided methods are trained from scratch and apply various techniques to reduce storage space. BBSplat achieved the highest storage space reduction compared to 3DGS (×$\times 17$ times) while providing state-of-the-art rendering quality.
@@
Qualitative comparison. We provide qualitative results in Figure 3 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art. For the DTU dataset, we also demonstrate better handling of challenging light effects. For more results, please refer to Supp. Mat..
@@
Mesh Extraction. In Table 4, we compare the 3D mesh extraction performance of BBSPlat with state-of-the-art methods on the DTU dataset [23]. For the experiment, we use 15 scenes with ground truth obtained by 3D scanning objects in a laboratory setting. For mesh extraction, we reduced input resolution to $800\times 600$ for compatibility with the values reported by 2DGS. As a metric, we present the Chamfer Distance (CD) indicating the similarity between two meshes. We also report PSNR for NVS with HD resolution and each 8th frame held for evaluation.
Both experiments for 2DGS and for BBSplat were conducted using the depth-normal regularization proposed in [20]. Rendering quality metrics for BBSplat in Tables 1 and 2 were also obtained using such regularizations. BBSplat reaches 0.91 CD on DTU dataset, which is better than 1.96 for 3DGS and 1.33 for SuGaR [17], and close to 0.80 for 2DGS. We provide extracted mesh visualization in Figure 4 (c). At the same time, BBSplat provides the best novel view rendering quality (PSNR) and requires the least storage space.
Ray-tracing rasterization. Representing the scene as textured billboards allows us to consider it as a set of planes forming a textured mesh. Thus, we can rasterize such a plane-based mesh using classical computer graphic rasterization, enabling ray-tracing effects as shown in Figure 4 (a). The image was obtained by rasterizing in Blender a set of planes with RGB and alpha textures. This way it is not necessary to spend extra time for extracting mesh representation and it allows to preserve photo realistic rendering while enabling rasterization effects and editing. For animated examples, please refer to the supplementary video.
To reduce the number of overlapping artifacts during rasterization, we used StopThePop [42] per-ray sorting algorithm during training. Note that StopThePop in this work was used only for visualizations in Figure 4. After training, we convert each billboard to the set of two polygons forming a plane with corresponding UV coordinates on the concatenated texture map.
[…]
3DGS [24] 1.96 28.39 121 MB
SuGaR [17] 1.33 27.95 1059 MB
2DGS [20] 0.80 28.04 59 MB
BBSplat(Ours) 0.91 29.72 32 MB